# Explainable Visual Question Answering with Multimodal Rationales

### v2

@@ IV-C2 Ablation of Unified or Separate Text Predictor
IV-C3 Ablation of the Text Generator
In our proposed approach, we use GPT-2 as the textual rationale generator. To evaluate its effectiveness compared to commonly used text generation models, such as LSTMs and Transformer, we replace this component in MRVQA-E with these alternative modules. The results are shown in Table VIII. Comparing the first three rows, GPT-2 significantly outperforms standard LSTMs and yields slightly better results than Transformer, even with frozen parameters. This comparison proves the similar observation on the effectiveness of the powerful pre-training in Table V. Subsequently, we allow the parameters of GPT-2 to be trainable within our MRVQA-E model to boost generation performance. While this requires additional effort to train the LLMs, it enhances the quality of the generated textual rationales. For instance, it results in a 13.78% improvement in the CIDEr score. This advancement in linguistic understanding also positively impacts visual predictions, as evidenced by a 5.82% increase in the vtS score. We further replace the text generator with the more advanced GPT-3 model [56], known for its superior performance in complex reasoning and contextual understanding. However, GPT-3’s size (175 billion parameters compared to GPT-2’s 1.5 billion) poses significant challenges, and moreover, it is not available for fine-tuning. Given our intention to fine-tune LLMs for easy adaptation and computational efficiency, we utilize GPT-3 with frozen parameters for generating textual rationales. The results indicate that using more advanced LLMs does not significantly impact the performance of text generation for our EVQA task.
TABLE VIII: Ablation results of different text generators under either freezing or fine-tuning settings on VQA-E-Syn.
@@ IV-C3 Ablation of the Text Generator
IV-C4 Ablation of Linguistic Features for Vision Predictor
To assess the impact of linguistic features on the vision predictor Grounding-DINO, we perform experiments on VQA-E-Syn using features derived from either questions ortextual rationales as input. We observe that the variant using features from input questions struggles (41.93% AP), as these do not provide comprehensive and clear referring information like visual grounding [57]. In contrast, features derived from textual rationales significantly enhance the detection model, leading to 47.45% AP (a 5.52% improvement) in the MRVQA-E model. In addition, Fig. 4 shows a qualitative comparison between the variant and our approach using both MRVQA-E and MRVQA-C models. The results demonstrate that our proposed approach leverages the generation of textual rationales to significantly enhance performance for multimodal EVQA.
Figure 4: Examples of generated visual rationales (bounding boxes in red) with different linguistic inputs for our MRVQA-C and MRVQA-E models.
IV-C5 Ablation of Loss Functions
To further evaluate the proposed loss function, we conduct ablation experiments on VQA-E-Syn to verify its effectiveness. The results are reported in Table IX. Comparing the first two rows, we can see that incorporating visual supervision significantly improves overall performance across all types of predictions, which is consistent with our earlier observations in Table II and Table III. Comparing the last four rows, we observe a notable enhancement in performance due to the inclusion of our proposed text alignment loss. For example, all configurations utilizing $Loss_{TA}$ show at least a 1% improvement in SPICE scores over the others. In addition, we perform a series of experiments to determine the optimal value for the trade-off parameter $\lambda$. The results from the last three rows indicate that setting $\lambda=0.5$ yields the best performance. This suggests that the balance component is relatively less important than the basic loss components for achieving the multimodal predictions.
TABLE IX: Ablation results of loss functions on VQA-E-Syn.
@@ IV-C5 Ablation of Loss Functions
[…]
IV-C6 Ablation of the New Metric
To verify whether the introduced metric vtS effectively measures the quality of the generated multimodal rationales, we perform a series of analyses in this part. Table X reports the results. We begin by evaluating the textual similarity score (TS) derived from the GTE [44] model. The results in the first two columns show that the quality of textual rationales, as represented by TS, aligns well with the SPICE metric. Additionally, we explore various combinations of visual and textual aspects in the multimodal rationale evaluation. Specifically, we consider three options: (1) averaging AP and TS by taking their mean; (2) multiplying AP and TS; (3) dividing the result from (2) by the result from (1), which We refer to as vtS. We observe that options (1) and (2) fail to accurately reflect model performance when there is a significant discrepancy between the two components (e.g., comparing VCIN and MRVQA-C, and considering that visual predictions would deteriorate further if the model’s performance were to decline). In contrast, option (3) mitigates the impact of large discrepancies, offering a more balanced assessment (e.g., 0.3% with (3) compared to more than 0.7% with (1)). Consequently, we use the combination approach outlined in (3) as the new evaluation metric vtS for our multimodal EVQA task.
TABLE X: Ablation results of the multimodal evaluation metric on VQA-E-Syn.
Method S TS AP (1) (2) (3)
PJ-X [8] 15.32 68.32 38.43 53.38 26.26 49.19
VQA-E [9] 16.83 71.67 40.65 56.16 29.13 51.88
FME [10] 18.77 73.16 42.57 57.87 31.14 53.82
DMRFNet [11] 20.41 75.06 44.74 59.90 33.58 56.06
VCIN [31] 22.07 78.13 45.97 62.05 35.92 57.89
MRVQA-C 22.19 79.57 45.86 62.72 36.49 58.19
MRVQA-E 23.68 78.56 47.45 63.00 37.28 59.16
@@ IV-C6 Ablation of the New Metric
Figure 5: Qualitative results of our approach compared to the previous methods. The bounding boxes in green display the relevant visual rationales.
IV-D Qualitative Analysis
@@
We conduct a qualitative analysis of predicted answers and rationales. Fig. 5 displays three examples from the synthesized multimodal dataset VQA-E-Syn. Compared to the previous methods such as DMRFNet [11] and VCIN [31], our proposed approach demonstrates superior performance in both answering and rationale support. For example, in the first column, both DMRFNet and VCIN struggle to comprehend the query’s intent, misleading them to focus on the object “woman,” although VCIN ultimately predicts the correct answer “No”. In contrast, our MRVQA model effectively performs cross-modal understanding, accurately identifying the target “man” referring to the statue and correctly predicting the answer, accompanied by reasonable rationales. Another example in the second column illustrates our model’s ability to capture relevant objects in both visual and textual rationales. Our MRVQA model successfully identifies the action “sitting on” and accurately locates the target, while the other methods are misled by “the other cat” or “the toilet”. Additionally, the final example demonstrates our EVQA model’s reasoning capability. Despite the lack of category-level information, the model accurately predicts the answer and delivers consistent multimodal rationales, as evidenced by the references to “the man” and “the tennis racket”. These examples demonstrate that our approach not only delivers accurate answers but also offers robust and reliable multimodal rationales, effectively explaining the prediction process.
V Discussion and Limitations
While our approach has shown advancements through extensive experiments across various datasets, there remain several areas for further exploration in future research. In our proposed model, we utilize the pre-trained CLIP model to represent input questions and images. However, the model may encounter problems when users’ focus is on local details within images or when contextual information is limited. We show two failure cases in Fig. 6 to illustrate these problems. In the left example, our model predicts the wrong answer and fails to identify the relevant objects due to the challenge of distinguishing the target among numerous similar objects. This issue may stem from the CLIP model’s strength in capturing global features from large-scale image-caption datasets, which might limit its ability to provide detailed local representations. In the example shown right, although our model accurately understands the cross-modal input and successfully locates the relevant objects, the limited contextual information hampers the precision of the linguistic prediction. The challenge arises because the zoomed-in local view complicates the recognition of the object (e.g., mirror). This scenario emphasizes the importance of a multimodal EVQA model, which provides clear insights compared to a “black box” answering system. Therefore, a promising direction for future work is to explore the usage of models with more advanced feature representations and enhanced cross-modal recognition capabilities, such as VilBert[24] and BLIP [4].
Figure 6: Two failure cases of our method.
Moreover, Multimodal Large Language Models (MLLMs) offer a superior approach for cross-modality tasks by leveraging mutual learning. Due to their significantly larger parameter sizes and extensive training datasets, these MLLMs often outperform domain-specific methods. They have gained attention for various vision-language tasks, owing to the capability to deliver high performance through downstream fine-tuning [54, 58]. However, these MLLMs usually respond slowly and demand substantial computational resources for training and fine-tuning. In addition, these models typically generate complex text based on the provided context (e.g., prompt), which may require users to understand or simplify the responses to obtain concise and straightforward answers. In our original intention for this multimodal EVQA task, we try to avoid redundant fine-tuning across multiple predictions while seeking a clean and computation-friendly solution for non-expert users. Recent advancements, such as adaptors (e.g., Lora [59]), significantly reduce the number of trainable parameters in MLLMs. This development facilitates further progress in EVQA research.
VI Conclusion
@@
In this paper, we propose a multimodal EVQA method that predicts not only answers but also textual and visual rationales to support the answers. To minimize the effort required for annotating precise visual labels, we synthesize a dataset by leveraging existing datasets through a semi-automatic matching and refinement engine. To solve the multimodal EVQA problem, we introduce MRVQA, a Transformer-based framework that integrates a projection module to align pre-trained features with large language models for generating textual rationales. Additionally, MRVQA employs an open-vocabulary detector to extract precise bounding boxes for consistent visual rationales. Furthermore, to the best of our knowledge, we are the first to propose a metric for evaluating the quality of generated multimodal explanations from a visual perspective. The extensive experiments across four VQA datasets demonstrate that our approach significantly outperforms existing state-of-the-art EVQA methods. We believe these works will inspire further research in the EVQA field.

### v3

@@ 4.3.2 Ablation of Unified or Separate Text Predictor
4.3.3 Ablation of the Text Generator
To evaluate LLMs’ effectiveness compared to commonly used text generation models, such as LSTMs and Transformer, we replaced this component in MRVQA-E with these alternative modules. The results are shown in Table 6. Comparing the first three rows, GPT-2 significantly outperformed standard LSTMs and yielded slightly better results than Transformer, even with frozen parameters. Subsequently, we allowed the parameters of GPT-2 to be trainable to boost generation performance. While this required additional effort to train the LLMs, it enhanced the quality of the generated textual rationales. For instance, it resulted in a 13.78% improvement in the CIDEr score. This advancement in linguistic understanding also positively impacted visual predictions, as evidenced by a 5.82% increase in the vtS score. We further replaced the text generator with the more advanced GPT-3 model [6], known for its superior performance in complex reasoning and contextual understanding. However, GPT-3’s size (175 billion parameters compared to GPT-2’s 1.5 billion) poses significant challenges, and moreover, it is not available for fine-tuning. We utilized GPT-3 with frozen parameters for generating textual rationales. The results indicate that using more advanced LLMs does not significantly impact the performance of text generation for our EVQA task.
Table 6: Ablation results of different text generators under either freezing or fine-tuning settings on VQA-E-Syn.
@@ 4.3.3 Ablation of the Text Generator
4.3.4 Ablation of Loss Functions
To further evaluate the proposed loss function, we conducted ablation experiments on VQA-E-Syn to verify its effectiveness. The results are reported in Table 7. Comparing the first two rows, we can see that incorporating visual supervision significantly improved overall performance across all types of predictions, which is consistent with our earlier observations in Table 1 and Table 2. Comparing the last four rows, we observed a notable enhancement in performance due to the inclusion of our proposed text alignment loss. For example, all configurations utilizing $Loss_{TA}$ showed at least a 1% improvement in SPICE scores over the others. In addition, we performed a series of experiments to determine the optimal value for the trade-off parameter $\lambda$. The results from the last three rows indicate that setting $\lambda=0.5$ yielded the best performance.
Table 7: Ablation results of loss functions on VQA-E-Syn.
@@ 4.3.4 Ablation of Loss Functions
15.84 22.19 - 71.65
15.13 22.34 59.01 72.28
$\lambda=0.1$ 16.49 23.51 59.10 74.76
$\lambda=0.5$ 16.97 23.68 59.16 75.01
$\lambda=1$ 16.25 23.40 59.07 74.59
@@
Figure 3: Qualitative comparison of results. The bounding boxes in green display the relevant visual rationales.
4.4 Qualitative Analysis
@@
Figure 3 displays three examples from the synthesized multimodal dataset VQA-E-Syn. Compared to the previous methods such as DMRFNet [50] and VCIN [47], our proposed approach demonstrates superior performance in both answering and rationale support. For example, in the first column, both DMRFNet and VCIN struggled to comprehend the query’s intent, focusing on the object “woman,” although VCIN ultimately predicted the correct answer “No”. In contrast, MRVQA effectively performed cross-modal understanding, accurately identifying the target “man” referring to the statue and correctly predicting the answer. The medium column illustrates our model’s ability to capture relevant objects in both visual and textual rationales. It successfully identified the action “sitting on” and located the target, while the other methods were misled by “the other cat” or “the toilet”. In the last example, despite the lack of category-level information, MRVQA accurately predicted the answer and delivered consistent rationales, as evidenced by the references to “the man” and “the tennis racket”.
5 Conclusion
@@
In this paper, we propose a multimodal EVQA method that predicts not only answers but also textual and visual rationales to support the answers. To minimize the annotation effort, we synthesize a dataset by augmenting existing datasets through a semi-automatic matching engine. To solve the multimodal EVQA problem, we introduce MRVQA that generates textual rationales with large language models and visual rationales via an pre-trained detector. A loss function is introduced to ensure the consistency between the answer and textual rationale generation. Furthermore, we propose a metric for evaluating the quality of generated rationales from a visual perspective. The extensive experiments across three EVQA datasets demonstrate that MRVQA significantly outperforms existing state-of-the-art EVQA methods. We believe this work will inspire further research in the EVQA field.
Acknowledgement Kun Li has been supported by China Scholarship Council. Michael Ying Yang has been supported by EU HORIZON-CL42023-HUMAN-01-CNECT XTREME (grant no.101136006).